a wire is bent in the form of a quadrant of a circle of diameter 1.4m. find the length of the wire, if the quadrant is bounded by the radii of the circle.
Answer by edjones(8007)

The radii are 0.7 each and the length of the arc is 1/4 of the circumference.
pi*D/4=pi*1.4/4=1.1
1.1+.7+.7=2.2 m
Ed